OpenStudy (anonymous):

Given f(x,y)=x^2+4y^2. using x=1 and y=1 and differentials, find an approximation for (1.01)^2 + 4(1.99)^2

OpenStudy (anonymous):

First get the Partial of x and y. So Fx =2x Fy = 8y

OpenStudy (anonymous):

Fx(1,1) = 2 Fy(1,1) = 8 f(1,1) = 5

OpenStudy (anonymous):

\[T(x,y) = Fx(x _{0},y _{0})\Delta x + Fy(x _{0},y _{0})\Delta y+ f(x _{0},y _{0})\]

OpenStudy (anonymous):

T= 2(x-1) + 8(y-1) + 5

OpenStudy (anonymous):

T= 2x-2 + 8y - 8 + 5 T= 2x +8y -5

OpenStudy (anonymous):

T((1.01^2, 4(1.99)^2) = 2((1.01)^2) + 8(4(1.99)^2) -5

OpenStudy (anonymous):

\[\approx 123.76\]

OpenStudy (anonymous):

If that isnt the correct answer I know it is atleast the right steps.